OpenStudy (anonymous):

The probability that Janice will attend math class today is 62%. The probability that Ellie will attend math class if Janice attends is 100%. Ellie will not attend the class unless she hears from Janice first. What is the probability that both will attend the class? 73% 100% 81% 62%

OpenStudy (anonymous):

Conditional probability: \[ \Pr(A|B)=\frac{\Pr(A\cap B)}{\Pr(B)} \]

OpenStudy (anonymous):

Which means that: \[ \Pr(A\cap B) = \Pr(A|B)\Pr(B) \]

OpenStudy (anonymous):

This makes no scene @wio

OpenStudy (anonymous):

You are not familiar with conditional probability?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

One event is that Janice attends. The other event is that Ellie attends.

OpenStudy (anonymous):

A condition event would be that Ellie attends if Janice attends.

OpenStudy (anonymous):

So...?

OpenStudy (anonymous):

So I think you might want to read your textbook a bit, since you're completely clueless.

OpenStudy (anonymous):

I do online school so i dont have a text book with me. Im doing a post test

OpenStudy (valpey):

Here are the possibilities: Janice attends and Ellie attends Janice attends and Ellie skips Janice skips and Ellie attends Janice skips and Ellie skips let A = "Janice attends" let B = "Ellie attends" We can rewrite this: P(A)*P(B|A) P(A)*P(~B|A) P(~A)*P(B|~A) P(~A)*P(~B|~A)

OpenStudy (valpey):

You would read that first line "the probability of A times the probability of B given A happens (is true)" You would read the third line "the probability of not A times the probability of B given not A (A is not true)"

OpenStudy (valpey):

Anyway, our probabilities are given in the problem: P(A)*P(B|A) = 62%*100% P(A)*P(~B|A) = 62%*0% P(~A)*P(B|~A) = 38%*0% P(~A)*P(~B|~A) = 38%*100%